Find a solution, if any, for the following system of equations: 421 – 5x2 + 1x3 = 4 -6x1 + 6x2+ = -6 3x1 + +2x3 = 8
Answer #1

solution: The given system of equations is 421 - 582 & 193=4 - 6x + 6x2 = -6 3x +ang = 8 Here the coefficient matrix is 6 6.A - 4 -6 6 3 8 det A x3 = 4648-0) +5(8+18) + 4(6-18) -30 system Therefore e the solution of the given of équation is x=2, 82=like please.